OpenStudy (anonymous):

if y varies inversely with x, and if y= 10 when x= 1.2, then what is the value of x when y=30 ? A) 0.4 B) 0.6 C) 0.8 D) 10 E) 12 Can u guys explain please :D

OpenStudy (akshay_budhkar):

do you know what inversley varying means?

OpenStudy (lalaly):

y=k/x 10=k/1.2 k=12 30=12/x x=12/30 = 0.4

OpenStudy (akshay_budhkar):

it is when y=k/x where k is a constant..

OpenStudy (akshay_budhkar):

oh!!! lana has done the job!! thanks lana!

OpenStudy (anonymous):

thank you guy :DDD